You are testing the claim that the proportion of men who own cats is larger than the proportion of women who own cats. Assume the null hypothesis is that the population proportions are equal.

You sample 160 men, and 30% own cats.

You sample 120 women, and 45% own cats.

Find the test statistic, rounded to two decimal places.
Answer #1

p = (p1 * n1 + p2 * n2) / (n1 + n2) = (0.3 * 160 + 0.45 * 120) / (160 + 120) = 0.364

SE = sqrt{ p * ( 1 - p ) * [ (1/n1) + (1/n2) ] } = sqrt(0.364 * (1 - 0.364) * (1/160 + 1/120)) = 0.0581

Test statistic = z = (p1 - p2) / SE = (0.3 - 0.45) / 0.0581 = -2.58